please i really need help on this mixture problem! i have tried but couldn't get the starting equation. here is the question:a banker has 10 pound bags full of coins. one bag is 70% pennies and the other is 30% pennies. how many pounds from each bag must the banker mix to get a 15 pound bag with 40% pennies?
Answer by stanbon(75887)

a banker has 10 pound bags full of coins. one bag is 70% pennies and the other is 30% pennies. how many pounds from each bag must the banker mix to get a 15 pound bag with 40% pennies?
-----
pennies + pennies = pennies
0.70*x + 0.30*(15-x) = 0.40*15
-----------------------------------
70x + 30*15 - 30x = 40*15
40x = 10*15
x = (1/4)15 = 3 3/4 lbs (amt of pennies from the 70% bag)
15-x = 11.25 lbs (amt. of pennies from the 30% bag)
-----
Cheers,
Stan H.
============